a 1.75 kg stone falls from the top of a cliff and strikes the ground at a speed of 58.8 m/s. what is the height of the cliff?
1
Expert's answer
2020-01-30T09:43:24-0500
"h=\frac{v^2}{2g}"

"h=\frac{58.8^2}{2(9.8)}=176\ m"